OpenStudy (htmerritt):

this question was on here 29 days ago, i believe the answer was wrong. It is missing 2J. Can someone please tell me where 2J in 2nd line went Four friends are collecting canned goods for charity. solving for r = ron jasmine = j barry = b alicia = a r = 3j + 10 b = 2j a = b + 10 -- a = 2j + 10 (j + r) = (a + b) (j + (3j + 10) = (2j + 10) + 2j 4j + 10 = 2j + 12 4j - 2j = 12 - 10 2j = 2 j = 1 r = 3j + 10 r = 3(1) + 10 r = 3 + 10 r = 13 b = 2j b = 2(1) b = 2 a = 2j + 10 a = 2(1) + 10 a = 2 + 10 a = 12 check... j + r = a + b 1 + 13 = 12 + 2 14 = 14 (correct)

OpenStudy (htmerritt):

Four friends are collecting canned goods for charity. Ron collects 10 more than 3 times as many cans as Jasmine. Barry collects twice as many cans as Jasmine. Alicia collects 10 more cans than Barry. How many cans does each person collect if Jasmine and Ron collect the same number of cans as Alicia and Barry?

OpenStudy (shinalcantara):

are you given the total goods?

OpenStudy (htmerritt):

no

OpenStudy (htmerritt):

ron and jasmine same as alicia and barry...I keep getting 4J+10=4J+10 which I think would be 0

OpenStudy (shinalcantara):

yeah it will be zero for sure since you're not given the total amount.. if that's the case there are numerous answers to the problem..

OpenStudy (shinalcantara):

From the given, you will have let 'x' be Ron's number of canned goods collected 'y' be Jasmine's 'z' be Barry's 'w' be Alicia's as stated: 'Ron collects 10 more than three times as many as Jasmine' x = 10 + 3y -----equation 1 Barry is having twice as Jasmine z = 2y ------equation 2 Alicia is having 10 more than Barry w = 10 + 2y ----equation 3 Ron's plus Jasmine's is equal to Barry's plus Alicia's x + y = z + w -----equation 4 You have 4 equations, 4 unknowns then you can solve for the values: plugging in equations 1,2 and 3 to equation 4 x + y = z + w 10 + 3y + y = 2y + 10 + 2y 4y + 10 = 4y + 10 NOTICE that the two expressions left and right are equal then you will have an infinite number of solutions -so y can be any value as long as it is not negative since you can't expect the number of canned goods to be negative right? and it should be in whole number for you can't expect a 1/3 or 1/4 number of canned goods ;)